James and Lin got married and got new jobs in Chicago. Lin earns $3400 more per year than James. Together they earn $82,300. How much do each of them earn per year?
Answer by wilft1(217)

x + x + 3400 = 82300
subtract 3400 from both sides
- 3400 - 3400
x + x = 78900
combine like terms
2x = 78900
divide both sides by 2
James is x = $39450
Lin is x + 3400 = $42850